OpenStudy (anonymous):

simplify x^2 - 16x + 64 over x^2 - 64

OpenStudy (anonymous):

x^2 - 16x + 64 = (x-8)^2 x^2 - 64 = (x+8)(x-8) now try it

OpenStudy (anonymous):

-1?

OpenStudy (anonymous):

\[\frac{(x-8)^2}{(x-8)(x+8)} = \frac{x-8}{x+8}\]